# Exact black hole solutions in gravity with a background Kalb-Ramond field

### v1

@@ VI CONCLUSIONS
In this work, we studied black hole solutions in the presence of a nonzero VEV of Kalb-Ramond field in the three-dimensional stationary axisymmetric spacetime and the four-dimensional static spherically symmetric spacetime and the three-dimensional stationary axisymmetric spacetime. We presented exact solutions for four-dimensional static spherically symmetric black holes and three-dimensional rotating black holes.
First, we noted that in some previous works, the variation of the nonminimal coupling term between the Kalb-Ramond field and the scalar curvature was neglected. Instead, this term was absorbed into a redefinition of variables. We have properly taken this term into account and obtained new four-dimensional spherically symmetric black hole solutions. These solutions modify the previous results given in Ref. Yang et al. (2023)., where the cosmological constant was $\Lambda=0$ for the potential $V(X)=\frac{\lambda}{2}X^{2}$, and $\Lambda\neq 0$ for $V(X)=\frac{\lambda}{2}X$. Additionally, by incorporating this term, we have also obtained new black hole solutions in the case of a nonzero cosmological constant for the potential $V(X)=\frac{\lambda}{2}X^{2}$. This allows for a direct consideration of black hole solutions with and without a cosmological constant under the potential $V(X)=\frac{\lambda}{2}X^{2}$. Due to Lorentz symmetry breaking in Kalb-Ramond gravity, the asymptotic behavior of the resulting spacetime Is neither Minkowskian nor (A)dS.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2020YFC2201503), the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, No. 12475055 and No. 12205129), the 111 Project (Grant No. B20063), Gansu Province’s Top Leading Talent Support Plan.

### v2

@@ VI CONCLUSIONS
In this work, we studied black hole solutions in the presence of a nonzero VEV of Kalb-Ramond field in the three-dimensional stationary axisymmetric spacetime and the four-dimensional static spherically symmetric spacetime. We presented exact solutions for four-dimensional static spherically symmetric black holes and three-dimensional rotating black holes.
First, we noted that in some previous works, the variation of the nonminimal coupling term between the Kalb-Ramond field and the scalar curvature was neglected. Instead, this term was absorbed into a redefinition of variables. We have properly taken this term into account and obtained new four-dimensional spherically symmetric black hole solutions. These solutions modify the previous results given in Ref. Yang et al. (2023), where the cosmological constant was $\Lambda=0$ for the potential $V(X)=\frac{\lambda}{2}X^{2}$, and $\Lambda\neq 0$ for $V(X)=\frac{\lambda}{2}X$. Additionally, by incorporating this term, we have also obtained new black hole solutions in the case of a nonzero cosmological constant for the potential $V(X)=\frac{\lambda}{2}X^{2}$. This allows for a direct consideration of black hole solutions with and without a cosmological constant under the potential $V(X)=\frac{\lambda}{2}X^{2}$.
@@
In summary, we have obtained new classes of black hole solutions in gravity with a background Kalb–Ramond field and analyzed their thermodynamic properties using the Iyer–Wald formalism. These solutions, characterized by Lorentz-violating parameters, introduce small but potentially observable deviations from general relativity. In the absence of a cosmological constant, the solution exhibits slight departures from asymptotic flat spacetime. These solutions provide a theoretical framework to probe Lorentz-symmetry violations in strong-gravity regimes. In particular, possible observational signatures—such as quasinormal modes, black hole shadows, and extreme mass-ratio inspirals—may offer promising avenues to test the physical relevance of these effects.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2020YFC2201503), the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, and No. 12475055 ), the 111 Project (Grant No. B20063), Gansu Province’s Top Leading Talent Support Plan.